# Some Super-approximation Rates of ReLU Neural Networks for Korobov Functions

### v1

@@ 2.2 Some Lemmas
Lemma 1 (Proposition 4.3 from Lu et al. (2021)).
@@
for $k=0,1,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},K-1$.
The following lemmas (cf. Siegel (2023); Jiao et al. (2023)) are useful for the analysis of neural networks.
Lemma 2.
@@
Lemma 3 (Proposition 4.4 from Lu et al. (2021)).
Let $W,L,s\in\mathbb{N}_{+}$ and $\xi_{i}\in[0,1]$ for $i=0,1,\ldots,W^{2}L^{2}-1$, there exists a neural network
@@
$\sum_{i=1}^{N_{1}N_{2}}\phi_{i}=\sum_{i=1}^{N_{2}}\sum_{j=1}^{N_{1}}\phi_{i,j},$
where each $\phi_{i,j}$ is still of width $W$ and depth $L$, then for each $i=1,2\ldots,N_{2}$, we have $\sum_{j=1}^{N_{1}}\phi_{i,j}\in\mathcal{NN}_{d,1}(N_{1}W,L)$ based on the first result in the fourth part of Lemma 2. By using the second result in the fourth part of Lemma 2 with $n=N_{2}$ and $k=1$ there, we further have $\phi\in\mathcal{NN}_{d,1}(N_{1}W+2d+2,N_{2}L)$. ∎
Lemma 5 (Proposition 4 from Yang et al. (2023b)).
@@
Next we construct a neural network that efficiently approximates the product of finite input variables in the $W^{1}_{\infty}$ norm over a larger bounded domain. Furthermore, we give the upper bound of the $W^{1}_{\infty}$ norm of the resulting network function. The proof of Lemma 6 is similar to (Hon and Yang, 2022, Lemma 3.5) with the help of Lemma 5. When the intervals $[-2,2]$ and $[-c,c]$ below are replaced by $[0,1]$, the result described corresponds to (Yang et al., 2023b, Proposition 5).
Lemma 6.
@@
$\left\|\phi(\bm{x})-x_{1}x_{2}\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu}x_{d+1}\right\|_{W^{1}_{\infty}\left(\Omega_{c}\right)}\leq 14a^{4}(W+1)^{-7dL}.$
Proof.
By Lemma 5, there exists a function $\phi_{1}\in\mathcal{NN}(15(W+1),14dL)$ such that $\|\phi_{1}\|_{W^{1}_{\infty}([-2^{d+1},2^{d+1}]^{2})}\leq 12\cdot 2^{2d+4}$ and
$\|\phi_{1}(x,y)-xy\|_{W^{1}_{\infty}([-2^{d+1},2^{d+1}]^{2})}\leq 6\cdot 2^{2d+4}\cdot(W+1)^{-7dL}.$ (1)
@@
$\Big{\|}\phi_{i}(x_{1},x_{2},\ldots,x_{i+1})-\prod_{j=1}^{i+1}x_{j}\Big{\|}_{W^{1}_{\infty}([-2,2]^{i+1})}$ (2)
$\leq$ $(2^{i}-1)\cdot 6\cdot 2^{2d+4}\cdot(W+1)^{-7dL},$
@@
By Lemma 2, $\phi_{i+1}\in\mathcal{NN}(15(W+1)+2i+1,14(i+1)dL)$. Using the error bounds (1) and (2) implies that
$\Big{\|}\phi_{i+1}(x_{1},x_{2},\ldots,x_{i+2})-\prod_{j=1}^{i+2}x_{j}\Big{\|}_{W^{1}_{\infty}([-2,2]^{i+2})}$
$\leq$ $(2^{i+1}+1)\cdot 6\cdot 2^{2d+4}(W+1)^{-7dL}.$
@@
$\|\phi(\bm{x})-x_{1}x_{2}\ldots x_{d}\|_{W^{1}_{\infty}([-2,2]^{d})}\leq 2^{3d+6}(W+1)^{-7dL}.$ (3)
Using Lemma 5 again with $a=\max\{2^{d+1},c\}$ , there is a function $\psi\in\mathcal{NN}(15(W+1),14dL)$ such that $\|\psi\|_{W^{1}_{\infty}\left([-a,a]^{2}\right)}\leq 12a^{2}$ and
$\|\psi(x,y)-xy\|_{W^{1}_{\infty}\left([-a,a]^{2}\right)}\leq 6a^{2}(W+1)^{-7dL}$ (4)
Let $\Phi(\bm{x}):=\psi(\phi(\bm{x}^{\prime}),x_{d+1})$ for any $\bm{x}=(\bm{x}^{\prime},x_{d+1})\in\mathbb{R}^{d+1}$. By Lemma 2, (3), and (4), we have $\Phi\in\mathcal{NN}(15(W+1)+2d-1,14d^{2}L)$, $\|\Phi\|_{W^{1}_{\infty}\left(\Omega_{c}\right)}\leq 12a^{2}$ and
$\left\|\Phi(\bm{x})-x_{1}x_{2}\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu}x_{d+1}\right\|_{W^{1}_{\infty}\left(\Omega_{c}\right)}\leq 14a^{4}\cdot(W+1)^{-7dL}.$
The proof is complete. ∎
Lemma 7.
@@
Proof.
We proceed by induction. For $m=1$, it is valid from the assumption. Assume the inequality holds for $m=k-1$, i.e.,
@@
By the principle of induction, the inequality holds for any $m\in\mathbb{N}_{+}$, thus we finish the proof. ∎
@@ 2.3 Interpolation on Sparse Grids
For any $f\in X^{2}_{p}(\Omega)$ with $|f|_{2,p}\leq 1$ and $n\in\mathbb{N}_{+}$, $1\leq p<\infty$, consider the lowest-order sparse grid interpolation (Bungartz and Griebel (2004))
$\Pi_{n}f(\bm{x})=\sum_{|\bm{l}|_{1}\leq n+d-1}\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}v_{\bm{l},\bm{i}}\phi_{\bm{l},\bm{i}}(\bm{x}),$ (5)
@@
$\mathcal{I}_{\bm{l}}:=\left\{\bm{i}\in\mathbb{N}_{+}^{d}:\bm{1}\leq\bm{i}\leq 2^{\bm{l}}-\mathbf{1},~{}i_{j}\text{ is odd for all }1\leq j\leq d\right\},$ (6)
and the coefficient $v_{\bm{l},\bm{i}}$ satisfying $|v_{\bm{l},\bm{i}}|\leq 1$. It is proved in Mao and Zhou (2022) that for a given $n$ satisfying
@@
$\|f-\Pi_{n}f\|_{L_{p}(\Omega)}\leq W^{-2}(\log W)^{3(d-1)}\leq 2^{-2n}(nd)^{3(d-1)}.$ (7)
Let $m\geq 3$, $1\leq p<\infty$, and $n\in\mathbb{N}_{+}$, for any $f\in X^{m}_{p}(\Omega)$ with $|f|_{m,p}\leq 1$, consider the $(m-1)$-th order sparse grid interpolation
$\Pi_{n}^{m}f(\bm{x})=\sum_{|\bm{l}|_{1}\leq n+d-1}\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}v_{\bm{l},\bm{i}}^{m}\phi_{\bm{l},\bm{i}}^{m}(\bm{x}),$ (8)
where $\phi_{\bm{l},\bm{i}}^{m}$ is the higher order sparse grid basis function of degree $\leq m-1$ in each direction (see Bungartz and Griebel (2004) and Li and Zhang (2025)), and the coefficient $v_{\bm{l},\bm{i}}^{m}$ satisfies $|v_{\bm{l},\bm{i}}^{m}|\leq C_{m,d}$ for some fixed constant $C_{m,d}$. For $n\in\mathbb{N}_{+}$, the sparse grid interpolation $\Pi_{n}^{m}f$ satisfies
$\|f-\Pi_{n}^{m}f\|_{L_{p}(\Omega)}\leq C_{m,d}2^{-mn}n^{d-1},$ (9)
@@ 3 Approximation by ReLU DNN in $L_{p}$ Norm
This section is devoted to the proof of Theorem 1. First we define a region for sparse grids by excluding a small area similar to that in Lu et al. (2021), which will be used in the proofs of Theorem 1 and Theorem 2.
Definition 2.
@@ 3.1.1 The case $m=2$
For any fixed multi-index $\bm{l}=(l_{1},\ldots,l_{d})\in\mathbb{N}_{+}^{d}$, applying Lemma 1 (with appropriate adjustments made to the parameters of the final affine transformation) with $K=2^{l_{j}-1}\leq W^{2}(2L)^{2}$ yields a function $\phi_{l_{j}}\in\mathcal{NN}(4W+3,8L+5)$ for each dimension $j=1,\ldots,d$, satisfying the following piecewise constant property:
$\phi_{l_{j}}(x)=2k-1\quad\text{ if }x\in\left[\frac{2k-2}{2^{l_{j}}},\frac{2k}{2^{l_{j}}}-\varepsilon\cdot\mathbf{1}_{\{k<2^{l_{j}}-1\}}\right],$ (10)
@@
By Lemma 2, we have $\Phi_{\bm{l}}^{1}\in\mathcal{NN}(4dW+3d,8L+5)$. Moreover, by (10), it holds that
$\Phi_{\bm{l}}^{1}(\bm{x})=\bm{i},\quad\text{ for any}\,\,\bm{x}\in\Omega_{\bm{i},\varepsilon}^{\bm{l}},\,\,\text{and for each}\,\,\bm{i}\in\mathcal{I}_{\bm{l}}.$ (11)
Let $h(\bm{x})=\frac{\bm{x}-1}{2}$ for any $\bm{x}\in\mathbb{R}^{d}$, then $h$ is a bijection from $\mathcal{I}_{\bm{l}}$ to $\prod_{j=1}^{d}\{0,1,\ldots,2^{l_{j}-1}-1\}$. In addition, for each $k\in\big{\{}0,1,\ldots,\prod_{j=1}^{d}2^{l_{j}-1}-1\big{\}}$, we could define a bijection
@@
$\Phi_{\bm{l}}:=\mathcal{A}_{1}\circ\Phi_{\bm{l}}^{1}\in\mathcal{NN}(4dW+3d,8L+5)$ (12)
@@
$\Phi_{\bm{l}}(\bm{x})=\operatorname{ind}(\bm{i}):$ $=\sum_{j=1}^{d}\frac{i_{j}-1}{2}\prod_{r=1}^{j}2^{l_{r}-1}$ (13)
$\in\left\{0,1,\ldots,\prod_{j=1}^{d}2^{l_{j}-1}-1\right\}.$
@@
$g_{1}\binom{\bm{z}_{1}}{\bm{z}_{2}}=\left(\Phi_{\bm{l}}^{1}(\bm{z}_{1}),\sigma(\bm{z}_{2})+\sigma(-\bm{z}_{2})\right)^{\top},$ (15)
for any $\bm{z}_{1},\bm{z}_{2}\in\mathbb{R}^{d}$. By Lemma 2 and (11), the composed network $g_{1}\circ\mathcal{A}_{2}$ satisfies
$g_{1}\circ\mathcal{A}_{2}\in\mathcal{NN}(4dW+5d,8L+5).$ (16)
Denote by $\bm{I}^{(i)}\in\mathbb{R}^{3\times d}$ the matrix where the element in column $i$ is $1$, while the elements in other columns are all 00 and $\bm{H}^{(i)}\in\mathbb{R}^{d\times 3}$ the matrix where the $i$-th row vector is $(1,-2,1)$, while the elements in other rows are all 00, for each $i=1,2,\ldots,d$. Let $\mathcal{A}_{3}(\bm{x}_{1})=\bm{W}_{3}\bm{x}_{1}+\bm{b}_{3}$ and $\mathcal{A}_{4}(\bm{x}_{2})=\bm{W}_{4}\bm{x}_{2}$ with
$\bm{W}_{3}=\begin{pmatrix}-\bm{I}^{(1)}&2^{l_{1}}\bm{I}^{(1)}\\ \vdots&\vdots\\ -\bm{I}^{(d)}&2^{l_{d}}\bm{I}^{(d)}\\ \bm{O}&\bm{I}_{d}\end{pmatrix}\in\mathbb{R}^{4d\times 2d},$
$\bm{W}_{4}=\begin{pmatrix}\bm{H}^{(1)}&\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu}&\bm{H}^{(d)}&\bm{O}\\ \bm{O}&\ldots&\bm{O}&\bm{I}_{d}\\ \end{pmatrix}\in\mathbb{R}^{2d\times 4d},$
and
$\bm{b}_{3}=(-1,0,1,-1,0,1,\ldots,-1,0,1,\bm{0}_{d}^{\top})^{\top}\in\mathbb{R}^{4d}$
for any $\bm{x}_{1}\in\mathbb{R}^{2d},\bm{x}_{2}\in\mathbb{R}^{4d}$, where $\bm{I}_{d}\in\mathbb{R}^{d\times d}$ is the identity matrix and $\bm{0}_{d}=(0,0,\ldots,0)^{\top}\in\mathbb{R}^{d}$. It follows from (16) that
$\phi_{1}:=\mathcal{A}_{4}\circ\sigma\circ\mathcal{A}_{3}\circ g_{1}\circ\mathcal{A}_{2}\in\mathcal{NN}(4dW+5d,8L+6).$ (17)
@@
$g_{2}\binom{\bm{z}_{1}}{\bm{z}_{2}}=\left(\sigma(\bm{z}_{1})+\sigma(-\bm{z}_{1}),\Phi_{\bm{l}}(\bm{z}_{2})\right)^{\top},$ (18)
for any $\bm{z}_{1},\bm{z}_{2}\in\mathbb{R}^{d}$. By Lemma 2 and equations (12) and (17), we obtain
$\widetilde{\phi}_{1}:=g_{2}\circ\phi_{1}\in\mathcal{NN}(4dW+7d,16L+11)$ (19)
@@
by Lemma 3, there exists a function
@@
$g_{3}\binom{\bm{z}_{1}}{z_{2}}=\left(\sigma(\bm{z}_{1})+\sigma(-\bm{z}_{1}),\widetilde{\phi}_{2}(z_{2})\right)^{\top},$
for any $\bm{z}_{1}\in\mathbb{R}^{d},z_{2}\in\mathbb{R}$. Using Lemma 2, we can deduce that
@@
ϕ3::subscriptitalic-$\phi_{3}:$ $=g_{3}\circ\widetilde{\phi}_{1}$
$\in\mathcal{NN}(16ds(W+1)\log(8W)+7d,(26L+21)\log(8L))$
@@
$|\widetilde{v}_{\bm{l},\bm{i}}-v_{\bm{l},\bm{i}}|\leq 2W^{-2s}(2L)^{-2s},\quad\text{for any}\,\,\bm{i}\in\mathcal{I}_{\bm{l}}.$ (20)
By Lemma 6, there exists a function
@@
Finally, let $\phi_{\bm{l}}:=\phi_{4}\circ\phi_{3}$; see Figure 1 for illustration. Using Lemma 2, we have
[…]
$(26L+21)\log(8L)+14(d+1)^{2}L).$
@@
for any $\bm{x}\in\Omega_{\bm{i},\varepsilon}^{\bm{l}}$ and for each $\bm{i}\in\mathcal{I}_{\bm{l}}$. Combining this result with (20) and $\|\phi_{\bm{l},\bm{i}}\|_{L_{\infty}(\Omega)}\leq 1$ implies that
@@
where $\Omega_{\varepsilon}:=\cap_{|\bm{l}|_{1}\leq n+d-1}\Omega_{\varepsilon}^{\bm{l}}$ and $\mu$ is the Lebesgue measure of a set in $\mathbb{R}^{d}$. In fact, we can calculate $\|f_{\bm{l}}\|_{L_{\infty}(\Omega)}$ (see Bungartz and Griebel (2004)) and $\|\phi_{\bm{l}}\|_{L_{\infty}(\Omega)}$ using Lemma 2.2 in Shen et al. (2019) for each $\bm{l}$ satisfying $|\bm{l}|_{1}\leq n+d-1$. Notably, both these norms only depend on $n$. Consequently, by selecting a sufficiently small $\varepsilon$, we can ensure the following holds
$\mu((\Omega\backslash\Omega_{\varepsilon}))(\|f_{\bm{l}}\|_{L_{\infty}(\Omega)}+\|\phi_{\bm{l}}\|_{L_{\infty}(\Omega)})$
$\leq$ $3^{5d+4}(W+1)^{-14(d+1)L}+2W^{-2s+2}(2L)^{-2s+2}.$
@@
Then, utilizing Lemma 4, (21) and the fact
$\sum_{|\bm{l}|_{1}\leq n+d-1}1$ =|$=|\{\bm{l}\in\mathbb{N}_{+}^{d}:|\bm{l}|_{1}\leq n+d-1\}|$ (23)
$\leq(n+d)^{d}\leq(2d)^{d}(\log(8W))^{d}(\log(8L))^{d},$
where the first inequality is a classical result in the literature (see also (Li and Li, 2025, Lemma 4.7)), the size of the network function satisfies

### v2

@@ 2.2 Some Lemmas
Lemma 2.1 (Proposition 4.3 from [26]).
@@
for $k=0,1,\cdots,K-1$.
The following lemmas (cf. [35, 18]) are useful for the analysis of neural networks.
Lemma 2.2.
@@
Lemma 2.3 (Proposition 4.4 from [26]).
Let $W,L,s\in\mathbb{N}_{+}$ and $0\leq\xi_{i}\leq 1$ for $i=0,1,\ldots,W^{2}L^{2}-1$, there exists a neural network
@@
$\phi=\sum_{i=1}^{N_{1}N_{2}}\phi_{i}=\sum_{i=1}^{N_{2}}\sum_{j=1}^{N_{1}}\phi_{i,j},$
where each $\phi_{i,j}$ is still of width $W$ and depth $L$, then for each $i=1,2\ldots,N_{2}$, we have $\sum_{j=1}^{N_{1}}\phi_{i,j}\in\mathcal{NN}_{d,1}(N_{1}W,L)$ based on the first result in the fourth part of Lemma 2.2. By using the second result in the fourth part of Lemma 2.2 with $n=N_{2}$ and $k=1$ there, we further have $\phi\in\mathcal{NN}_{d,1}(N_{1}W+2d+2,N_{2}L)$.
Lemma 2.6 (Proposition 4 from [38]).
@@
Next we construct a neural network that efficiently approximates the product of finite input variables in the $W^{1}_{\infty}$ norm over a larger bounded domain. Furthermore, we give the upper bound of the $W^{1}_{\infty}$ norm of the resulting network function. The proof of Lemma 2.7 is similar to [17, Lemma 3.5] with the help of Lemma 2.6. When the intervals [2,2][-2,2] and [c,c][-c,c] below are replaced by $[0,1]$, the result described corresponds to [38, Proposition 5].
Lemma 2.7.
@@
$\left\|\phi(\bm{x})-x_{1}x_{2}\cdots x_{d+1}\right\|_{W^{1}_{\infty}\left(\Omega_{c}\right)}\leq 14a^{4}(W+1)^{-7dL}.$
Proof 2.8.
By Lemma 2.6, there exists a function $\phi_{1}\in\mathcal{NN}(15(W+1),14dL)$ such that $\|\phi_{1}\|_{W^{1}_{\infty}([-2^{d+1},2^{d+1}]^{2})}\leq 12\cdot 2^{2d+2}$ and
$\|\phi_{1}(x,y)-xy\|_{W^{1}_{\infty}([-2^{d+1},2^{d+1}]^{2})}\leq 6\cdot 2^{2d+2}\cdot(W+1)^{-7dL}.$ (2.1)
@@
$\Big\|\phi_{i}(x_{1},x_{2},\ldots,x_{i+1})-\prod_{j=1}^{i+1}x_{j}\Big\|_{W^{1}_{\infty}([-2,2]^{i+1})}\leq(2^{i}-1)\cdot 6\cdot 2^{2d+2}\cdot(W+1)^{-7dL},$ (2.2)
@@
By Lemma 2.2, $\phi_{i+1}\in\mathcal{NN}(15(W+1)+2i+1,14(i+1)dL)$. Using the error bounds (2.1) and (2.2) implies that
$\Big\|\phi_{i+1}(x_{1},x_{2},\ldots,x_{i+2})-\prod_{j=1}^{i+2}x_{j}\Big\|_{W^{1}_{\infty}([-2,2]^{i+2})}\leq(2^{i+1}-1)\cdot 6\cdot 2^{2d+2}(W+1)^{-7dL}.$
@@
$\|\phi(\bm{x})-x_{1}x_{2}\ldots x_{d}\|_{W^{1}_{\infty}([-2,2]^{d})}\leq 2^{3d+6}(W+1)^{-7dL}.$ (2.3)
Using Lemma 2.6 again with $a=\max\{2^{d+1},c\}$ , there is a function $\psi\in\mathcal{NN}(15(W+1),14dL)$ such that $\|\psi\|_{W^{1}_{\infty}\left([-a,a]^{2}\right)}\leq 12a^{2}$ and
$\|\psi(x,y)-xy\|_{W^{1}_{\infty}\left([-a,a]^{2}\right)}\leq 6a^{2}(W+1)^{-7dL}$ (2.4)
Let $\Phi(\bm{x}):=\psi(\phi(\bm{x}^{\prime}),x_{d+1})$ for any $\bm{x}=(\bm{x}^{\prime},x_{d+1})\in\mathbb{R}^{d+1}$. By Lemma 2.2, (2.3), and (2.4), we have $\Phi\in\mathcal{NN}(15(W+1)+2d-1,14d^{2}L)$, $\|\Phi\|_{W^{1}_{\infty}\left(\Omega_{c}\right)}\leq 12a^{2}$ and
$\left\|\Phi(\bm{x})-x_{1}x_{2}\cdots x_{d+1}\right\|_{W^{1}_{\infty}\left(\Omega_{c}\right)}\leq 14a^{4}\cdot(W+1)^{-7dL}.$
The proof is complete.
Lemma 2.9.
@@
Proof 2.10.
We proceed by induction. For $m=1$, it is valid from the assumption. Assume the inequality holds for m=k1m=k-1, i.e.,
@@
By the principle of induction, the inequality holds for any $m\in\mathbb{N}_{+}$, thus we finish the proof.
@@ 2.3 Interpolation on Sparse Grids
For any $f\in X^{2}_{p}(\Omega)$ with $|f|_{2,p}\leq 1$ and $n\in\mathbb{N}_{+}$, $1\leq p<\infty$, consider the lowest-order sparse grid interpolation ([9])
$\Pi_{n}f(\bm{x})=\sum_{|\bm{l}|_{1}\leq n+d-1}\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}v_{\bm{l},\bm{i}}\phi_{\bm{l},\bm{i}}(\bm{x}),$ (2.5)
@@
$\mathcal{I}_{\bm{l}}:=\left\{\bm{i}\in\mathbb{N}_{+}^{d}:\bm{1}\leq\bm{i}\leq 2^{\bm{l}}-\mathbf{1},~i_{j}\text{ is odd for all }1\leq j\leq d\right\},$ (2.6)
and the coefficient $v_{\bm{l},\bm{i}}$ satisfying $|v_{\bm{l},\bm{i}}|\leq 1$. It is proved in [27] that for a given $n$ satisfying
@@
$\|f-\Pi_{n}f\|_{L_{p}(\Omega)}\leq W^{-2}(\log W)^{3(d-1)}\leq 2^{-2n}(nd)^{3(d-1)}.$ (2.7)
Let $m\geq 3$, $1\leq p<\infty$, and $n\in\mathbb{N}_{+}$, for any $f\in X^{m}_{p}(\Omega)$ with $|f|_{m,p}\leq 1$, consider the (m1)(m-1)-th order sparse grid interpolation
$\Pi_{n}^{m}f(\bm{x})=\sum_{|\bm{l}|_{1}\leq n+d-1}\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}v_{\bm{l},\bm{i}}^{m}\phi_{\bm{l},\bm{i}}^{m}(\bm{x}),$ (2.8)
where $\phi_{\bm{l},\bm{i}}^{m}$ is the higher order sparse grid basis function of degree $\leq m-1$ in each direction (see [9] and [22]), and the coefficient $v_{\bm{l},\bm{i}}^{m}$ satisfies $|v_{\bm{l},\bm{i}}^{m}|\leq C_{m,d}$ for some fixed constant $C_{m,d}$. For $n\in\mathbb{N}_{+}$, the sparse grid interpolation $\Pi_{n}^{m}f$ satisfies
$\|f-\Pi_{n}^{m}f\|_{L_{p}(\Omega)}\leq C_{m,d}2^{-mn}n^{d-1},$ (2.9)
@@ 3 Approximation by ReLU DNN in $L_{p}$ Norm
This section is devoted to the proof of Theorem 1.2. First we define a region for sparse grids by excluding a small area similar to that in [26], which will be used in the proofs of Theorem 1.2 and Theorem 1.3.
Definition 3.1.
@@ 3.1.1 The case $m=2$
For any fixed multi-index $\bm{l}=(l_{1},\ldots,l_{d})\in\mathbb{N}_{+}^{d}$, applying Lemma 2.1 (with appropriate adjustments made to the parameters of the final affine transformation) with $K=2^{l_{j}-1}\leq W^{2}(2L)^{2}$ yields a function $\phi_{l_{j}}\in\mathcal{NN}(4W+3,8L+5)$ for each dimension $j=1,\ldots,d$, satisfying the following piecewise constant property:
$\phi_{l_{j}}(x)=2k-1\quad\text{ if }x\in\left[\frac{2k-2}{2^{l_{j}}},\frac{2k}{2^{l_{j}}}-\varepsilon\cdot\mathbf{1}_{\{k<2^{l_{j}}-1\}}\right],$ (3.10)
@@
By Lemma 2.2, we have $\Phi_{\bm{l}}^{1}\in\mathcal{NN}(4dW+3d,8L+5)$. Moreover, by (3.10), it holds that
$\Phi_{\bm{l}}^{1}(\bm{x})=\bm{i},\quad\text{ for any}\,\,\bm{x}\in\Omega_{\bm{i},\varepsilon}^{\bm{l}},\,\,\text{and for each}\,\,\bm{i}\in\mathcal{I}_{\bm{l}}.$ (3.11)
Let $h(\bm{x})=\frac{\bm{x}-\bm{1}}{2}$ for any $\bm{x}\in\mathbb{R}^{d}$, then $h$ is a bijection from $\mathcal{I}_{\bm{l}}$ to $\prod_{j=1}^{d}\{0,1,\ldots,2^{l_{j}-1}-1\}$. In addition, for each $k\in\big\{0,1,\ldots,\prod_{j=1}^{d}2^{l_{j}-1}-1\big\}$, we could define a bijection
@@
$\Phi_{\bm{l}}:=\mathcal{A}_{1}\circ\Phi_{\bm{l}}^{1}\in\mathcal{NN}(4dW+3d,8L+5)$ (3.12)
@@
$\Phi_{\bm{l}}(\bm{x})=\operatorname{ind}(\bm{i}):=\sum_{j=1}^{d}\frac{i_{j}-1}{2}\prod_{r=1}^{j}2^{l_{r}-1}\in\left\{0,1,\ldots,\prod_{j=1}^{d}2^{l_{j}-1}-1\right\}.$ (3.13)
@@
$g_{1}\binom{\bm{z}_{1}}{\bm{z}_{2}}=\left(\Phi_{\bm{l}}^{1}(\bm{z}_{1}),\sigma(\bm{z}_{2})+\sigma(-\bm{z}_{2})\right)^{\top},$ (3.15)
for any $\bm{z}_{1},\bm{z}_{2}\in\mathbb{R}^{d}$. By Lemma 2.2 and (3.11), the composed network $g_{1}\circ\mathcal{A}_{2}$ satisfies
$g_{1}\circ\mathcal{A}_{2}\in\mathcal{NN}(4dW+5d,8L+5).$ (3.16)
Denote by $\bm{I}^{(i)}\in\mathbb{R}^{3\times d}$ the matrix where the element in column $i$ is $1$, while the elements in other columns are all 0 and $\bm{H}^{(i)}\in\mathbb{R}^{d\times 3}$ the matrix where the $i$-th row vector is (1,2,1)(1,-2,1), while the elements in other rows are all 0, for each $i=1,2,\ldots,d$. Let $\mathcal{A}_{3}(\bm{x}_{1})=\bm{W}_{3}\bm{x}_{1}+\bm{b}_{3}$ and $\mathcal{A}_{4}(\bm{x}_{2})=\bm{W}_{4}\bm{x}_{2}$ with
$\bm{W}_{3}=\begin{pmatrix}-\bm{I}^{(1)}&2^{l_{1}}\bm{I}^{(1)}\\ \vdots&\vdots\\ -\bm{I}^{(d)}&2^{l_{d}}\bm{I}^{(d)}\\ \bm{O}&\bm{I}_{d}\end{pmatrix}\in\mathbb{R}^{4d\times 2d},\,\,\bm{W}_{4}=\begin{pmatrix}\bm{H}^{(1)}&\cdots&\bm{H}^{(d)}&\bm{O}\\ \bm{O}&\ldots&\bm{O}&\bm{I}_{d}\\ \end{pmatrix}\in\mathbb{R}^{2d\times 4d},$
and $\bm{b}_{3}=(-1,0,1,-1,0,1,\ldots,-1,0,1,\bm{0}_{d}^{\top})^{\top}\in\mathbb{R}^{4d}$ for any $\bm{x}_{1}\in\mathbb{R}^{2d},\bm{x}_{2}\in\mathbb{R}^{4d}$, where $\bm{I}_{d}\in\mathbb{R}^{d\times d}$ is the identity matrix and $\bm{0}_{d}=(0,0,\ldots,0)^{\top}\in\mathbb{R}^{d}$. It follows from (3.16) that
$\phi_{1}:=\mathcal{A}_{4}\circ\sigma\circ\mathcal{A}_{3}\circ g_{1}\circ\mathcal{A}_{2}\in\mathcal{NN}(4dW+5d,8L+6).$ (3.17)
@@
$g_{2}\binom{\bm{z}_{1}}{\bm{z}_{2}}=\left(\sigma(\bm{z}_{1})-\sigma(-\bm{z}_{1}),\Phi_{\bm{l}}(\bm{z}_{2})\right)^{\top},$ (3.18)
for any $\bm{z}_{1},\bm{z}_{2}\in\mathbb{R}^{d}$. By Lemma 2.2 and equations (3.12) and (3.17), we obtain
$\widetilde{\phi}_{1}:=g_{2}\circ\phi_{1}\in\mathcal{NN}(4dW+7d,16L+11)$ (3.19)
@@
by Lemma 2.3, there exists a function
@@
$g_{3}\binom{\bm{z}_{1}}{z_{2}}=\left(\sigma(\bm{z}_{1})-\sigma(-\bm{z}_{1}),\widetilde{\phi}_{2}(z_{2})\right)^{\top},$
for any $\bm{z}_{1}\in\mathbb{R}^{d},z_{2}\in\mathbb{R}$. Using Lemma 2.2, we can deduce that
@@
$\phi_{3}:=g_{3}\circ\widetilde{\phi}_{1}\in\mathcal{NN}(16ds(W+1)\log(8W)+7d,(26L+21)\log(8L))$
@@
$|\widetilde{v}_{\bm{l},\bm{i}}-v_{\bm{l},\bm{i}}|\leq 2W^{-2s}(2L)^{-2s},\quad\text{for any}\,\,\bm{i}\in\mathcal{I}_{\bm{l}}.$ (3.20)
By Lemma 2.7, there exists a function
@@
Finally, let $\phi_{\bm{l}}:=\phi_{4}\circ\phi_{3}$; see Figure 1 for illustration. Using Lemma 2.2, we have
$\phi_{\bm{l}}\in\mathcal{NN}(16ds(W+1)\log(8W)+7d,(26L+21)\log(8L)+14(d+1)^{2}L).$
@@
for any $\bm{x}\in\Omega_{\bm{i},\varepsilon}^{\bm{l}}$ and for each $\bm{i}\in\mathcal{I}_{\bm{l}}$. Combining this result with (3.20) and $\|\phi_{\bm{l},\bm{i}}\|_{L_{\infty}(\Omega)}\leq 1$ implies that
@@
where $\Omega_{\varepsilon}:=\cap_{|\bm{l}|_{1}\leq n+d-1}\Omega_{\varepsilon}^{\bm{l}}$ and $\mu$ is the Lebesgue measure of a set in $\mathbb{R}^{d}$. In fact, we can calculate $\|f_{\bm{l}}\|_{L_{\infty}(\Omega)}$ (see [9]) and $\|\phi_{\bm{l}}\|_{L_{\infty}(\Omega)}$ using Lemma 2.2 in [30] for each $\bm{l}$ satisfying $|\bm{l}|_{1}\leq n+d-1$. Notably, both these norms only depend on $n$. Consequently, by selecting a sufficiently small $\varepsilon$, we can ensure the following holds
$\mu((\Omega\backslash\Omega_{\varepsilon}))^{\frac{1}{p}}(\|f_{\bm{l}}\|_{L_{\infty}(\Omega)}+\|\phi_{\bm{l}}\|_{L_{\infty}(\Omega)})\leq 3^{5d+4}(W+1)^{-14(d+1)L}+2W^{-2s+2}(2L)^{-2s+2}.$
@@
Then, utilizing Lemma 2.4, (3.21) and the fact
$\sum_{|\bm{l}|_{1}\leq n+d-1}1$ $=|\{\bm{l}\in\mathbb{N}_{+}^{d}:|\bm{l}|_{1}\leq n+d-1\}|\leq(n+d)^{d}\leq(2d)^{d}(\log(8W))^{d}(\log(8L))^{d},$ (3.23)
where the first inequality is a classical result in the literature (see also [19, Lemma 4.7]), the size of the network function satisfies